# Kiloparsec-scale turbulence driven by reionization may grow intergalactic magnetic fields

### v1

@@ IV.2 The turbulent dynamo & B-field growth
Turbulence will act to grow magnetic fields. If a fraction $f_{B}$ of the turbulent kinetic energy goes into magnetic fields near the time of reionization, this would result in present-day $B$ fields of
@@
where the physical coherence length is also dilated by $1+z_{\rm re}$ by expansion. There is substantial disagreement about the nature of magnetic fields generated by turbulence as summarized in Schekochihin [44]. While simulations generally suggest that the field saturates with $f_{B}\approx 1$, some have argued that the $B$-field from turbulence saturates on scales near the driving scale of turbulence, with the magnetic field spectrum tracing the turbulence [23]. Others have argued that the scale should be substantially smaller [22].
The tightest constraints on IGM magnetic fields come from TeV blazars. Ackermann et al. [32] finds that the lack of detection of extended point sources requires $B_{\rm IGM}>3\times 10^{-16}(\lambda/10~{}{\rm kpc})^{-1/2}$ for $B$-field coherence lengths of $\lambda\lesssim 10~{}$kpc in their conservative analysis where they assume each blazar has been active for only ten years; the constraints rule out 30×30$\times$30 × larger $B_{\rm IGM}$ if the blazar is assumed to have its current TeV spectrum for $10^{4}$ years. Similar constraints have also been placed from a lack of GeV reprocessed photons in the spectral energy distribution [30, 31]. If we evaluate equation (5) at $z_{\rm re}=6$ and $c_{s}=20~{}$km s-1, the conservative constraint in Ackermann et al. [32] implies the $B$-field must have coherence of at least
@@
in order to explain the blazar constraints, where $f_{V}$ is the volume filling fraction of the turbulence and, hence, magnetic fields.
@@
Almost all pictures for the turbulent dynamo suggest that reionization-driven turbulence would result in magnetic fields that satisfy the constraint given by equation (6). In the accepted picture for the initial development of (non-helical) turbulent dynamo, the magnetic field grows until it comes into near equipartition with the kinetic energy of the eddies on the viscous scale so that the magnetic energy is $\approx V^{2}\,{\rm Re}^{-1/2}$. At this point, the magnetic field stretching by the viscous eddies gets suppressed, leading to ‘Batchelor level’ saturation [45, 24]. The growth during this stage happens on the short turnover timescale of the smallest hydrodynamic eddies ($L/[V~{}{\rm Re}^{1/2}]$). However, the scale of magnetic fluctuations is concentrated at the resistive scale when this saturation first happens (the scale over which magnetic diffusion is faster than the turnover time), which is given by $k_{\eta}={\rm Pr}^{1/2}k_{\nu}$, and we find magnetic Prandtl numbers of ${\rm Pr}\sim 10^{15}$ for our fiducial inputs. If this were the end of the story, even volume-filling turbulence would result in present-day coherence lengths of $\lambda\sim 10^{5}$ km and $f_{B}\sim 10^{-2}$, comparable to what is required to explain the blazar constraint given by equation (6).
[…]
@@
When the turbulence is not volume filling, as we find occurs at $T_{\rm min}\gtrsim 100~{}$K, there is an additional condition that must be satisfied to deflect the pairs from TeV blazars and hence suppress their GeV halos. Namely, each $e^{\pm}$ pair should intersect over its path a region where there has been turbulent amplification of $B_{\rm IGM}$. At $z=0$, each pair travels $\approx 400(E_{e}/1~{}\text{TeV})^{1/2}$ kpc before losing its energy by Compton scattering the CMB, where $E_{e}$ is the energy of one of the pairs, which is approximately half the energy of the original seed photon. The typical energy of the up-scattered CMB photons is then $2~{}\text{GeV}(E_{e}/1~{}\text{TeV})^{2}$, and constraints that find missing energy are centered at $\approx 1~{}$GeV, extending in some cases to $\approx 10~{}$GeV [e.g. 31]. Such photons arise from $0.5-3~{}$TeV pairs and, hence, these particles travel $200-800~{}$kpc. The 250250250~{}250kpc$/h$ slices shown in Figure 2 illustrate that for $T_{\rm min}=100~{}$K, even a 200200200~{}200kpc path would likely intercept a turbulent domain in which magnetic fields had been amplified.
@@ V Conclusions
In this work, we have shown that the destruction of mini-halos and IGM filaments by reionization drives turbulent eddies in the IGM that may fill a significant fraction of the IGM volume at $z>4$. This turbulence has the potential to amplify magnetic fields in the IGM, possibly explaining lower limits on IGM B-field strengths from TeV blazars. Our main findings are summarized below:
@@
  • We find that the energy spectrum of the turbulent gas follows a Kolmogorov scaling at wavenumbers $200\lesssim k/[{\rm Mpc}^{-1}]\lesssim 1000$ at $z<6$ for gas reionized at $z_{\rm re}=7$, assuming our fiducial resolution. The expected scaling extends to smaller scales with increasing spatial resolution. At $k\approx 10^{3}$ Mpc-1, we find that $\approx 1$ eddy turnover time has elapsed by $z=4$ for patches of gas reionized at $z=7$.
  • We showed that such turbulence would likely generate magnetic fields that could explain the lack of GeV halos towards TeV blazars, at least for $X$-ray preheating temperatures of $T\lesssim 100~{}$K. This result we argued depends little on the current debate regarding the character of the magnetic fields that would be generated by the turbulent dynamo. Even in conservative turbulent dynamo scenarios, the fields would likely be sufficient to explain the lack of pair halos if path lengths of $\sim 500~{}$kpc are likely to intersect a turbulent domain.
If reionization-driven turbulence is the source of IGM magnetic fields, future $\gamma$-ray observations, utilizing also echoes and off-axis blazars, may be able to differentiate reionization-turbulence dynamo pictures [e.g. 48]. Detecting magnetic fields from reionization turbulence would in turn imply that the pre-reionization IGM was only preheated to $\approx 100~{}$K by the time of reionization and that the inflationary density fluctuations were present down to mass scales of $\sim 10^{5}M_{\odot}$.
@@ Acknowledgements.
The authors thank Simon Foreman for sharing his dedicated computational resources on the Sol supercomputer at Arizona State University. Some computations for this work were run using computational resources distributed under NSF ACCESS allocations TG-PHY230158 and TG-PHY240332 and by the Sol supercomputer at Arizona State University. MM acknowledges support from NSF grant AST-2007012 and NASA grant 80NSSC24K1220. ES acknowledges support from NASA Grant 80NSSC22K1265. AD acknowledges support from NSF grant AST-2045600. HT acknowledges support from NASA grant 80NSSC22K0821.
@@ A.1 Spatial resolution
The viscosity of diffuse, ionized hydrogen is very small [49, Reynolds number $\sim 10^{5}$ for kpc-scale driving at the sound speed], such that flows are likely to be turbulent if a driving force is present. Shortly after reionization, pressure imbalances are everywhere as the IGM relaxes, providing such a force. Here, we analytically estimate the minimum spatial resolution required to capture this behavior, and ask whether it is consistent with our findings. The maximum characteristic length scale of IGM turbulence is set by the distance, $\Delta x^{\rm turb}$, that gas can travel in the time since it was reionized, $\Delta t$. If gas travels at roughly the sound speed, $c_{\rm s}$, then this scale is
@@
The limiting factor in capturing turbulence is “numerical viscosity” set by the simulation resolution, $\Delta x_{\rm cell}$. Teyssier [38] estimated the associated “numerical Reynolds number”, ${\rm Re}_{\rm num},$ to be $\approx 2L/\Delta x$ for their grid-based hydrodynamics simulations, where $L$ is the length scale associated with the turbulence (their Eq. 35, see also Rieder and Teyssier [50]). Taking their estimate and replacing $L$ with $\Delta x^{\rm turb}$ yields
@@ A.2 IGM pre-heating requirement
where we have taken densities in units of the cosmic mean of $\Delta\approx 10$ as the typical (initial) density of a structure responsible for generating turbulence in the diffuse IGM. Eq. 12 is consistent with our finding that turbulence disappears for $T_{\min}=1000~{}$K (see supplementary material), since structures with $\Delta\gg 10$ are rare in our tiny volumes. It also explains why turbulence only forms around the most massive structures in our $T_{\min}=100~{}$K run, since the above condition becomes $T<150~{}$K for $\Delta=1$ gas. Eq. 12 confirms our numerical results and suggests that the emergence of small-scale turbulence should be sensitive to the level and timing of pre-heating before reionization.
Appendix B Sensitivity to the self-shielding treatment
We test here whether our findings are sensitive to the implementation of self-shielding in our simulations. Throughout this work, we use the same setup for sending I-fronts across RT domains described in Cain et al. [14]. We run the simulation with hydrodynamics and gravity only until the reionization redshift, $z_{\rm reion}$. At this time, we begin tracing rays from the boundaries of RT domains, and “freeze” both cosmic time and hydrodynamical evolution until the I-fronts have crossed the RT domains. This ensures that the gas throughout the box responds coherently to the effects of reionization, and that the volume is fully reionized at exactly $z_{\rm reion}$. During this time, the reduced speed of light is set to 5×5$\times$5 × the speed of the I-front crossing the RT domain, or its full value, whichever is smaller, to ensure the correct I-front speed and post-I-front gas temperature [2, 3]. At $z<z_{\rm reion}$, we use a reduced speed of light of $\tilde{c}=2\times 10^{-3}$, which we find does not affect the self-shielding properties of the gas significantly [^4] Because of the tiny light-crossing time at the cell size, these simulations become computationally intractable if the speed of light is not reduced significantly. . We show here that this procedure does not affect our qualitative results, but does simplify their interpretation.
We have run two additional simulations with $N=1024^{3}$. The first flash-ionizes the box at $z=z_{\rm reion}$ (just like our $N=2048^{3}$) run, and thus has no self-shielding. The second sends a single plane-parallel I-front across the entire box at $z=z_{\rm reion}$ from just one direction, and does not freeze the density field during I-front passage. Respectively, these two approaches minimize and maximize the effect of self-shielding on IGM gas dynamics.
Appendix C Test with Nyx
Figure 5: Comparison of IGM turbulence in RadHydro and Nyx. We show slices through the temperature map at $z=6$, $5$, and $4$ (left to right) for our fiducial RadHydro run, a Nyx run with matching resolution, and our highest resolution RadHydro run (top to bottom). We find qualitatively similar turbulence in Nyx to that observed in RadHydro. Indeed, Nyx shows more turbulence on smaller scales at a fixed resolution than RadHydro, suggesting its algorithm is less diffusive.
Appendix D Supplemental Material
D.1 Conditions for achieving turbulence
D.1.1 Spatial resolution

### v2

@@ IV The turbulent dynamo & B-field growth
If a fraction $f_{B}$ of the turbulent kinetic energy goes into magnetic fields near the time of reionization, this would result in present-day $B$ fields of
@@
where $m_{p}$ is the proton mass, $n_{e}$ the electron density, and $V_{\rm dr}$ is the velocity of the driving. We have assumed the present-day field is reduced from that generated by $(1+z_{\rm re})^{-2}$, as would occur from homogeneous cosmological expansion, and we note that the physical coherence length is also dilated by $1+z_{\rm re}$ by expansion. There is substantial disagreement about the nature of magnetic fields generated by turbulence as summarized in Schekochihin [53]. While simulations generally suggest that the field saturates with $f_{B}\sim 1$, some have argued that the $B$-field from turbulence saturates on scales near the driving scale of turbulence, with the magnetic field spectrum tracing the turbulence [29]. Others have argued that the scale should be much smaller [28].
The tightest constraints on IGM B-fields come from TeV blazars. Neronov and Vovk [33] and Ackermann et al. [38] find that the lack of detection of extended point sources requires $B_{\rm IGM}>3\times 10^{-16}(\lambda/10~{\rm kpc})^{-1/2}$G for $B$-field coherence lengths of $\lambda\lesssim 10$~kpc in their conservative analysis where they assume each blazar has been active for only ten years; the constraints rule out $30\times$ larger $B_{\rm IGM}$ if the blazar is assumed to have its current TeV spectrum for $10^{4}$ years. Similar constraints have also been placed from a lack of GeV reprocessed photons in the spectral energy distribution [36, 37]. If we evaluate equation (2) at $z_{\rm re}=6$ and $c_{s}=20$~km s-1, the conservative constraint in Ackermann et al. [38] implies the $B$-field must have coherence of at least
@@
to explain the blazar constraints, where $f_{V}$ is the volume filling fraction of the turbulence and, hence, magnetic fields.
@@
Models for turbulent dynamo suggest that reionization-driven turbulence would result in magnetic fields that satisfy the constraint given by equation (3). In the accepted picture for the initial development of (non-helical) turbulent dynamo, the magnetic field grows until it comes into near equipartition with the kinetic energy of the eddies on the viscous scale so that the magnetic energy is $\approx V^{2}\,{\rm Re}^{-1/2}$. At this point, the magnetic field stretching by the viscous eddies gets suppressed, leading to ‘Batchelor level’ saturation [54, 30]. The growth during this stage happens on the short turnover timescale of the smallest hydrodynamic eddies ($L/[V~{\rm Re}^{1/2}]$). However, the scale of magnetic fluctuations is concentrated at the resistive scale when this saturation first happens (the scale over which magnetic diffusion is faster than the turnover time), which is given by $k_{\eta}={\rm Pr}^{1/2}k_{\nu}$, where Pr is the magnetic Prandtl number, and we find ${\rm Pr}\sim 10^{15}$ for our fiducial inputs. If this were the end of the story, even volume-filling turbulence would result in present-day coherence lengths of $\lambda\sim 10^{5}$ km and $f_{B}\sim 10^{-2}$, comparable to what is required to explain the blazar constraint given by equation (3). This would be true even in cosmic voids far from star forming halos, of which our tiny boxes are likely representative.
Hydrodynamic driven-turbulence simulations show that magnetic fields grow to energies above the Batchelor level saturation (closer to equipartition with the kinetic energy at the driving scale) and achieve larger coherence lengths than $k_{\eta}^{-1}$ [30]. However, the applicable regime of $1\ll\text{Re}\ll\text{Pr}$, is challenging to simulate and so studies tend to combine both analytic arguments and numerical results to draw conclusions about the saturated state. Schekochihin [53] suggested that instabilities associated with the tearing of eddy-aligned structures result in the field’s power spectrum peaking at $({\rm Re\times Pr})^{1/3}/L$, or roughly a length-scale of $10^{10}$~km for our fiducial inputs. With some numerical support, Yousef et al. [55] and Schekochihin [53] conjectured that the magnetic energy power spectrum may scale to lower wavenumbers in a manner such that there is equal magnetic energy per logkk, and, given sufficient time, extend down to $k\approx L^{-1}$. Any such redistribution of energy to larger scales would further act to deflect the Blazar TeV pairs.
A common hypothesis that is supported by some simulations [56] is that turbulence saturates with the magnetic field having the same energy distribution as the turbulent kinetic energy [e.g. 30, 53]. Xu and Lazarian [29] developed a quantitative theory for how such saturation would occur. Using their equations, we find that if there is $\approx 1$ turnover time at the driving scale as our simulations suggest (Appendix B), the field would have cascaded to match the turbulent kinetic energy down to $k\approx 10~L^{-1}$, which would be easily sufficient to quench the blazar pair halos. We note, however, that precisely evaluating the growth of the magnetic field in such a scenario would require MHD simulations, since magnetic fields with energy comparable to the kinetic energy of the gas can modify the turbulent dynamics at any scale, if not non-ideal MHD simulations that resolve down to the kilometer scales over which magnetic fields diffuse.
@@
When the turbulence is not volume filling, as we find occurs at $T_{\rm min}\gtrsim 100$~K, there is an additional condition that must be satisfied to deflect the pairs from TeV blazars and hence suppress their GeV halos. Namely, each $e^{\pm}$ pair should intersect over its path a region where there has been turbulent amplification of $B_{\rm IGM}$. At $z=0$, each pair travels $\approx 400(E_{e}/1~\text{TeV})^{1/2}$ kpc before losing its energy by Compton scattering the CMB, where $E_{e}$ is the energy of one of the pairs. The typical energy of the up-scattered CMB photons is then $2~\text{GeV}(E_{e}/1~\text{TeV})^{2}$, and constraints that find missing energy are centered at $\approx 1$~GeV, extending in some cases to $\approx 10$~GeV [e.g. 37]. Such photons arise from 0.530.5-3~TeV pairs and, hence, these particles travel 200800200-800~kpc. The 250250~kpc/h/h slices shown in Figure 2 illustrate that for $T_{\rm min}=100$~K, even a 200200~kpc path would likely intercept a turbulent domain in which magnetic fields had been amplified. The existence of turbulence in the $T_{\rm min}=100$ K case suggests that most regions in the IGM would still see some turbulence in this scenario.
@@ V Conclusions
For the first time, we have shown that the destruction of mini-halos and IGM filaments by reionization drives turbulent eddies in the IGM that may fill a significant fraction of the IGM volume at $z>4$. Our main findings are summarized below:
@@
  • We showed that such turbulence would likely generate magnetic fields that could explain the lack of GeV halos towards TeV blazars, at least for $X$-ray preheating temperatures of $T\lesssim 100$~K. This conclusion is likely robust to current debates regarding the character of the magnetic fields that would be generated by turbulence. Reionization driving turbulence represents a plausible mechanism to amplify volume-filling magnetic fields throughout the IGM.
If reionization-driven turbulence is the source of IGM magnetic fields, future $\gamma$-ray observations, utilizing also echoes and off-axis blazars, may be able to differentiate reionization-turbulence dynamo pictures [e.g. 57].
@@ Acknowledgements.
The authors thank Simon Foreman for sharing his dedicated computational resources on the Sol supercomputer at Arizona State University. Some computations for this work were run using computational resources distributed under NSF ACCESS allocations TG-PHY230158 and TG-PHY240332 and by the Sol supercomputer at Arizona State University. CC acknowledges support from the Beus Center for Cosmic Foundations. MM acknowledges support from NSF grant AST-2007012 and NASA grant 80NSSC24K1220. ES acknowledges support from NASA Grant 80NSSC22K1265. AD acknowledges support from NSF grant AST-2045600. HT acknowledges support from NASA grant 80NSSC22K0821.
@@ A.1 Spatial resolution
The viscosity of diffuse, ionized hydrogen is very small [58, Reynolds number $\sim 10^{5}$ for kpc-scale driving at the sound speed], such that flows are likely to be turbulent if a driving force is present. Shortly after reionization, pressure imbalances are everywhere as the IGM relaxes, providing such a force. Here, we analytically estimate the minimum spatial resolution required to capture this behavior, and ask whether it is consistent with our findings. The maximum characteristic length scale of IGM turbulence is set by the distance, $\Delta x^{\rm turb}$, that gas can travel in the time since it was reionized, $\Delta t$. Assuming the gas speed in the turbulent eddies is comparable to the sound speed in ionized gas, $c_{s}$, then this scale is
@@
where $20$ km/s is taken as a typical reference value for $c_{s}$. The limiting factor in capturing turbulence is “numerical viscosity” set by the simulation resolution, $\Delta x_{\rm cell}$. Teyssier [48] estimated the associated “numerical Reynolds number”, ${\rm Re}_{\rm num},$ to be $\approx 2L/\Delta x$ for their grid-based hydrodynamics simulations, where $L$ is the length scale associated with the turbulence (their Eq. 35, see also Rieder and Teyssier [59]). Taking their estimate and replacing $L$ with $\Delta x^{\rm turb}$ yields
@@ A.2 IGM pre-heating requirement
where we have taken densities in units of the cosmic mean of $\Delta\approx 10$ as the typical (initial) density of a structure responsible for generating turbulence in the diffuse IGM. Eq. 9 is consistent with our finding that turbulence disappears for $T_{\min}=1000$~K (see supplementary material), since structures with $\Delta\gg 10$ are rare in our tiny volumes. It also explains why turbulence only forms around the most massive structures in our $T_{\min}=100$~K run, since the above condition becomes T<150T<150~K for $\Delta=1$ gas. Eq. 9 confirms our numerical results and suggests that the emergence of small-scale turbulence should be sensitive to the level and timing of pre-heating before reionization.
Appendix B Quantifying turbulence
In this appendix, we quantify the properties of turbulence in our simulations. The turnover timescale at wavenumber $k=2\pi/L$ in a turbulent flow is
$\tau_{\rm eddy}(k)\approx\frac{L}{\sigma_{\rm v}(k)}=\frac{2\pi/k}{[2kE(k)]^{1/2}},$ (10)
$E(k)\equiv\sum_{i}\Delta_{\rm v_{i}}^{2}(k)/(2k)$ (11)
is the so-called “energy spectrum”, where $\Delta_{\rm v_{i}}^{2}$ is the dimensionless power spectrum of the $i^{\rm th}$ velocity component. Many studies have shown that the Kolmogorov scaling of $E(k)\propto k^{-5/3}$ tends to be followed over the turbulent inertial range for isotropic, homogeneous turbulence [e.g. 61]. The total number of turnover times elapsed at scale $k$, $N_{\rm\tau}$, is given by integrating $\tau_{\rm eddy}^{-1}$ over time, starting from the time a sound wave can cross the distance scale $2\pi/k$ since the start of reionization.
Figure 3: Quantification of turbulence in our simulations. Top: The energy spectrum $E(k)$ at a time-series of redshifts, offset vertically for visualization, with redshift decreasing from bottom to top. The “bump” that moves to the left with time tracks the growing sound crossing distance since $z_{\rm re}$. At smaller scales, we find the $k^{-5/3}$ scaling characteristic of turbulence, with a fall-off at very small scales due to numerical dissipation. Bottom: the same as the top panel except without rescaling the amplitude and concentrating on $z=5$ and different spatial resolutions (gray curves). The high-$k$ fall off shifts to the right as resolution increases, as expected if this cutoff is numerical. The scaling at smaller $k$ is close to, but slightly shallower than, the 5/3-5/3 Kolmogorov scaling. The blue dotted line is a $0.5$ $h^{-1}$Mpc box with the fiducial resolution, and shows convergence of $E(k)$ with box size. The magenta-dashed curve shows a scenario with high X-ray pre-heating ($T_{\min}=1000$ K), for which we do not observe the 5/3-5/3 scaling.
In the top panel of Figure 3, we show $E(k)$ at a number of redshifts (decreasing from bottom to top) for our fiducial simulation. We use an arbitrary vertical scaling and offset the curves in sequence to easily compare their shapes. We offset the curves vertically so they can be easily compared, and show the expected $k^{-5/3}$ Kolmogorov scaling as a thin line. At $z=7.2$, just before ionization, $E(k)$ is slightly steeper than Kolmogorov, indicating that the dynamics are not dominated by turbulence. A small amount of turbulence may be present in the densest structures due to non-linear collapse [39], but if this is present it fills a small fraction of the volume. Just after ionization at $z=6.9$, $E(k)$ has a prominent “bump” centered at $\approx 800$ hh~Mpc-1. The right edge of the bump at $\approx 1500$ $h$ Mpc-1 corresponds to a co-moving distance scale of $\approx 4$ $h^{-1}$kpc, roughly twice the distance that a $20$ km s-1 sound wave can travel through a $T\approx 2\times 10^{4}$K highly photo-ionized gas in the $15$ Myr since $z_{\rm re}=7$. It is thus reasonable to interpret the bump as tracking the sound crossing distance since reionization. At successively lower redshifts, the bump shifts to smaller $k$ as sound waves travel further from their source, disappearing by $z\approx 5$ as the sound crossing distance approaches the box scale.
We expect turbulence to form at scales smaller than the sound-crossing distance – that is, to the right of the bump. This is exactly what we see – starting at $z=6.5$, $E(k)$ reaches the $k^{-5/3}$ scaling at a small range of $k$ around $600$ hhMpc-1. At lower redshifts, the $k$ range where this scaling holds increases, reaching as low as $\approx 200$ hhMpc-1 by $z=4$. The evolution of $E(k)$ with time quantitatively confirms our earlier assertion that the pressure-smoothing of small-scale structure is the driving mechanism of turbulence in our simulations, as opposed to structure formation [39] or some other mechanism. The snapshots before and close to $z_{\rm re}$ do not show a spectrum with a shape characteristic of turbulent motions, and this behavior only emerges $\sim 200$ Myr after $z_{\rm re}$. This time delay is consistent with the timescale required for expanding filaments to begin overlapping.
In the bottom panel, we show $E(k)$ at $z=5$ for simulations with varying spatial resolution (black/gray curves). The scaling at turbulent scales is slightly shallower than $k^{-5/3}$ – the green dot-dashed curve is best fit by a $k^{-1.45}$ scaling. This may be because turbulent eddies do not fill the entire volume, that the scale of turbulent driving $L\approx c_{s}\Delta t$ is increasing with time, and/or that some non-turbulent kinetic motion is present. The fall-off in $E(k)$ shifts to the right with increasing resolution, confirming that this is due to numerical dissipation. We also show $E(k)$ for a simulation with box size $L=0.5$ $h^{-1}$cMpc (blue dotted) and $T_{\rm min}=10^{3}$K (magenta dashed). The first demonstrates that the energy spectrum at turbulent scales is converged with respect to box size, and the second shows that extreme X-ray pre-heating erases turbulence (the magenta curve does not follow the $k^{-5/3}$ scaling), as qualitatively observed in Figure 2.
In reality, when the turbulence is hydrodynamic the cascade should extend to $k_{\nu}\approx{\rm Re}^{3/4}/L$ where the Reynolds number for mean density gas is given by
$\times\left(\frac{1+z}{7}\right)^{3}\left(\frac{T}{2\times 10^{4}{\rm K}}\right)^{-5/2},$
where $\nu$ is the kinematic viscosity for mean-density ionized gas. This indicates that the hydrodynamic turbulence can extend to sub-parsec scales for kpc driving and applicable parameters. As discussed in the next section, the short turnover time of these smaller eddies (with the smallest eddies turning over a time $\text{Re}^{-1/2}$ times shorter than the largest) allows turbulence to grow magnetic fields even though we find that the smallest eddies we can resolve turnover only $N_{\rm\tau}\approx 1$ time by $z=4$. We find that $N_{\tau}$ is smaller than this when pre-heating by X-rays is significant (as low as $\approx 0.1$ for $T_{\min}=1000$ K, see Supplemental Material for more detail).
Appendix C Supplemental Material
C.1 Conditions for achieving turbulence
C.1.1 Spatial resolution